OpenStudy (anonymous):

A recursive function is shown below: f(1) = 5 and f(n) = f(n - 1) - 6; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 5, 1, 7, 13, 19, ... 5, -1, -7, -13, -19, ... 5, 11, 17, 23, 29, ... 5, -11, -17, -23, -29, ... i really dont get these @iGreen.

OpenStudy (cwrw238):

work out the second term by putting n = 2 in the definition of f(n) what do you get?

OpenStudy (anonymous):

Aw, sorry I can't help you on these..I never got sequences either.. :(

OpenStudy (anonymous):

-5

OpenStudy (cwrw238):

f(n) = f(n-1) - 6 n = 2 f(2) = f(2-1) - 6 = f(1) - 6 = ?

OpenStudy (anonymous):

im so confused :(

OpenStudy (cwrw238):

what does the question say f(1) is?

OpenStudy (anonymous):

5

OpenStudy (cwrw238):

ok - so f(1) - 6 = 5-6

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

so the answer is B

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

f(3) = f(2) - 6 =-1-6 = -7 etc...

OpenStudy (cwrw238):

that's what recursive means - you keep on applyig the rulle - its like a loop in a computer program

OpenStudy (cwrw238):

each term is worked out from the previous term